Daytime aGw4 check - 20181003 UT


We were unable to locate the light source from the ARGOS cal unit. After much searching and debugging we believe that there are no photons coming out of the fiber at the center of the secondary. Gustavo is going to the mountain on Thursday, so he will have a look at that. See IT 7437.

In the meantime, there is nothing (beyond clouds) stopping you from trying AGW4 and LUCI1 for science. The first risk that the guide probe is not in precisely the correct place at the rotator center (which affects the precision of dithering). The second risk is that the wfs alignment has shifted slightly. I will come back to the remote room during twilight (if the weather clears) to look and see if there are any other WFS alignment issues. If it collimates, it's good to go.


James and Dan replaces the RAD motor/encoder unit in AGw4 yesterday evening, and Petr verified that the probe seems to be moving OK now. Now we are going to check that the probe is still correctly centered relative to the rotator axis.


20:00 UT Gustavo deploying SX ARGOS swing arm with assistance of Pat and Elliott. And he is having trouble with it - the park position switch seems not to be being activated as it should when arm is parked. See IT 7436.

20:16 Brake1 readback error on LFBG rotator. Elliott will reset the controller. That didn't help.

20:36 Dan R. went up to check the connectors - it recovered when Dan tightened the brake housing.

20:48 Guider has 10000 counts in 100 msec with binning=1.

21:10 Locking and unlocking the SX swing arms cured the problem that Gustavo was having, so now the ARGOS swing arm is deployed.

21:30 Doug bringing up LUCI1 to look for the source. We see a 2 arcmin circle, but seems not to be the ARGOS source.

22:19 Init and Home SX M2 and M3.

22:38 Air compressors have tripped off and panic'd primary mirror.

Pupil image: /Repository/20180903/luci1.20181004.0001.fits with SX ARGOS swing arm in.

22:45 Gustavo retracts SX ARGOS swing arm with assistance from Elliott.

Pupil image with ARGOS swing arm retracted 0003, 0004 and ARGOS source ON.

Pupil image with ARGOS swing arm retracted 0005 and ARGOS source OFF.

22:58 return telescope to Josh and Pat.

-- %USERSIG{JohnHill - 2018-10-03}%
Topic revision: r2 - 03 Oct 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback